OpenStudy (anonymous):

Is this correct? find f(-10) f(x)=-x - 8 f(-10)=-10-8 f=-18 I'm confuse

OpenStudy (precal):

f(-10)=-(-10)-8=10-8=2

OpenStudy (anonymous):

yes that true I forgot that the two negatives cancel each other.

OpenStudy (precal):

-(-) turn into a positive

OpenStudy (anonymous):

thank you

OpenStudy (precal):

anytime :)